A car moves with a speed of45km/h. Find the distance travelled by the car in 30 minutes

Grade:7

4 Answers

Arun
6 years ago
Dear Sourav
 
Speed = 45 km/hr
Means
In 1 hour or 60 kinutes it travels 45 km
Hence in 30 mins it will travel 45/2 = 22.5 Km
 
Regards
Arun (askIITians forum expert)
Ayan
5 years ago
Speed = 45 KM/HR.
1 hour = 60 mins = 45 Kms
30 mins= 45/2
= 22.5 or 22 ½ km per 30 min.
REENU SIVADARSHINI.M
5 years ago
speed =45km/ hr ; time =1/2  hr
 distance = speed * time
distance =45/2=22.5 km  
      so therefore the distance travelled is 22.5  km with 45 km /hr speed in ½ an  hour
Sayantan Garai
5 years ago
The car goes 45km in 1 hour or 60 min.
Hence it goes 45/2=22.5km in 60/2=30 min.
................................................................................